Solve the following pair of linear equation by matrix method 2p +q =5


5p +3=11

1
Expert's answer
2021-11-15T16:47:21-0500

The coefficient matrix is:


"A=\begin{pmatrix}
 2 & 1 \\
 5 & 0
\end{pmatrix}"

The variable matrix is:


"X=\begin{pmatrix}
 p \\
 q
\end{pmatrix}"

The constant matrix is:


"B=\begin{pmatrix}
 5 \\
 8
\end{pmatrix}"

Thus, to solve a system "AX=B," for "X," multiply both sides by the inverse of "A"


"A^{-1}AX=A^{-1}B"

and we shall obtain the solution:


"X=A^{-1}B"

Provided the inverse "A^{-1}" exists, this formula will solve the system.


"\det A=|A|=\begin{vmatrix}
 2 & 1 \\
 5 & 0
\end{vmatrix}=2(0)-1(5)=-5\not=0"

The inverse "A^{-1}" exists.


"A^{-1}=\dfrac{1}{-5}\begin{pmatrix}
 0 & -1 \\
 -5 & 2
\end{pmatrix}=\begin{pmatrix}
 0 & 1/5 \\
 1 & -2/5
\end{pmatrix}"

"X=A^{-1}B=\begin{pmatrix}
 0 & 1/5 \\
 1 & -2/5
\end{pmatrix}\begin{pmatrix}
 5 \\
 8
\end{pmatrix}"

"=\begin{pmatrix}
 0(5)+(1/5)(8) \\
 1(5)+(-2/5)(8)
\end{pmatrix}=\begin{pmatrix}
 8/5 \\
 9/5
\end{pmatrix}"

"\begin{pmatrix}
 p \\
 q
\end{pmatrix}=\begin{pmatrix}
 8/5 \\
 9/5
\end{pmatrix}"

"p=8\/5, q=9\/5"